OpenStudy (anonymous):

A Bus leaves a station and travels north at 75 kph. Two hours later a second bus leaves same station traveling north at 100 kph. How far from the station and after how many hours will the second bus overtake the first bus?

OpenStudy (anonymous):

How to solve this using two variables?

OpenStudy (compassionate):

You want to use a system of equations, hun.

OpenStudy (debbieg):

I would just let t=time since the FIRST bus left the station. Then at any time t: 1st bus' distance from the station =75t and 2nd bus' distance from the station =100(t-2) You want to find t when the distances are the same, so 75t=100(t-2) 200=25t t=8 Check that back in the original problem to make sure it works: After 8 hours, the first bus is 75*8=600 miles away from the station. The 2nd bus is 100*6=600 miles away.

OpenStudy (anonymous):

how do you solve it using x and y variables?

OpenStudy (anonymous):

you can put y instead of t. what matters?